# On the conditioning of histopolationL. Bruni Bruno, S. Serra-Capizzano

### v1

@@ Abstract
Histopolation is the approximation procedure that associates a degree d1d-1 polynomial $p_{d-1}\in\mathscr{P}_{d-1}(I)$ with a locally integrable function $f$ imposing that the integral (or, equivalently, the average) of $p$ coincides with that of $f$ on a collection of $d$ distinct segments $s_{i}$. In this work we discuss unisolvence and conditioning of the associated matrices, in an asymptotic linear algebra perspective, i.e., when the matrix-size $d$ tends to infinity. While the unisolvence is a rather sparse topic, the conditioning in the unisolvent setting has a uniform behavior: as for the case of standard Vandermonde matrix-sequences with real nodes, the conditioning is inherently exponential as a function of $d$ when the monomial basis is chosen. In contrast, for an appropriate selection of supports, the Chebyshev basis of second kind has a bounded conditioning. A linear behavior is also observed in the Frobenius norm.
@@
Polynomial histopolation; conditioning; Chebyshev polynomials.
{AMS}
15A12, 41A10, 42A15, 65D05
@@ 1 Introduction
Let $I\subset\mathbb{R}$ be an interval and let $f:I\to\mathbb{R}$ be a locally integrable function. Histopolation is the approximation procedure that associates with $f$ a degree d1d-1 polynomial $p_{d-1}\in\mathscr{P}_{d-1}(I)$ such that
(1) $\int_{\alpha_{i}}^{\beta_{i}}f(x)\mathrm{d}x=\int_{\alpha_{i}}^{\beta_{i}}p(x)\mathrm{d}x,$
for an appropriate collection of supports $s_{i}:=[\alpha_{i},\beta_{i}]$, $i=1,\ldots,d$. In the literature [5, 21], an equivalent formulation of (1) containing normalization by the measure $|s_{i}|$ of $s_{i}$ is sometimes adopted; in what follows, we stick with the non-normalized case. The main consequence of this choice is that the measures of the supports must be bounded away from zero; that is, we assume that there exists $\varepsilon>0$ independent of $d$ such that $|s_{i}|\geq\varepsilon$ for each $i=1,\ldots,d$. This is a rather natural assumption: even in the normalized case, it avoids the introduction of technical concepts such as regular sets [18]. We will also discuss limiting situations where segments collapse to points.
In contrast with the nodal interpolation procedure, where any collection of $d$ distinct points is unisolvent for polynomial interpolation in $\mathscr{P}_{d-1}(I)$, such a neat recipe does not exist for histopolation, where the identification of specific classes is a hard task. It is indeed very easy to construct examples of $d$ distinct segments that are not unisolvent for histopolation in $\mathscr{P}_{d-1}(I)$.
Example.
Consider the family of segments [i,i][-i,i], for $i=1,\ldots,d$ with $d\geq 1$. Any non-vanishing linear combination of polynomials of odd degree less or equal than d1d-1 satisfies
@@
This is only one of the many differences between nodal interpolation and histopolation. Such a dissimilarity is mostly motivated by the fact that any interval $s_{i}$ requires two parameters to be described (e.g., its endpoints $\alpha_{i}$ and $\beta_{i}$, or the midpoint $\tau_{i}$ and the length $\rho$), giving rise to a generally underdetermined problem. As a direct consequence of this, unisolvence results are confined to specific classes of segments. Many of them were identified in [4], and are:
@@
    Figure 1: The segments $s_{1}$, $s_{2}$ and $s_{3}$ form a chain of intervals. They are examples of segments in the class (C1).
@@
    segments with uniform arc-length. For some $\tau_{i}\in(0,\pi/2)$ and $0<\rho\leq\pi/2$, one has $\alpha_{i}=\cos(\tau_{i}+\rho)$ and $\beta_{i}=\cos(\tau_{i}-\rho)$;
@@
Of course, there are families of segments that belong to more than one class. An example of this are the so-called Chebyshev segments, that we will study in depth in Section 3. Indeed, for a particular choice of the arc-length parameter $\rho$, they can be framed both in class (C1) and class (C2). By exploiting the normalized version of (1), segments solving the Fekete problem have subsequently been recognized as an additional scenario [5]. In Section 2.3, we exhibit a new class of segments which is unisolvent for polynomial histopolation. We will denote this class by (C4).
@@
(2) $\left[V_{d}\right]_{i,j}:=\int_{s_{i}}b_{j}(x)\mathrm{d}x.$
As many of our results are asymptotic or directly related with the number of segments, denoted by $d$, we will consistently index the Vandermonde matrix $V_{d}$. This quantity is also the size of the matrix, and is upshifted by $1$ with respect to the relative polynomial degree.
Example.
The non-unisolvence of the set of segments introduced in the above example is also readily seen by looking at the Vandermonde matrix computed with respect the monomial basis $b_{j}(x)=x^{j-1}$. For any even $j$, one has
@@
It follows that any even row of $V_{d}$ is made out of zeros. Since the invertibility of the Vandermonde matrix does not depend on the polynomial basis chosen, one deduces that such a set is not unisolvent for histopolation in $\mathscr{P}_{d-1}(I)$.
The problem of computing the features of a Vandermonde matrix, in particular concerning its conditioning and hence the identification of ill-conditioned bases, is a cornerstone of approximation theory [20]; of course, so is also the opposite problem. In the framework of nodal interpolation, the problem was extensively studied, and several results concerning optimality [12, 13, 16], bad conditioning [2, 6], as well as their applications [17], have been obtained. It is worth pointing out that the majority of these results can be grouped into two clusters: results concerning the monomial basis, which in general offers unstable results, and results concerning the Chebyshev (of first kind) basis, which can be used in combination with Chebyshev nodes to reverse such a situation.
Outline of the paper and main results. In contrast with nodal interpolation, very few is known about the spectral features of the Vandermonde matrix (2) associated with the histopolation problem. The principal aim of this work is to fill this gap.
To do so, we dedicate the next section to ill-conditioning. We recognize that the (Euclidean) conditioning $\kappa_{2}$ of the matrix $V_{d}$ grows exponentially with the variable $d$ when the monomial basis is chosen. This result is presented in Section 2, first for the reference inverval I=[1,1]I=[-1,1] and then extended to any collection of segments by Theorem 2.13. This feature, which recalls the nodal case, is not a consequence of the mean value theorem. To broaden the range of applicability of the result, we also identify a new class of unisolvent segments, and use it to produce numerical evidences.
To restore hope in histopolation, in Section 3 we present results concerning well-conditioning. Indeed, we show that Chebyshev polynomials of the second kind, in combination with segments in the class (C2), give rise to a column-orthogonal matrix. The diagonal terms of this matrix are explicitly determined in Proposition 3.1. As a consequence, in Theorem 3.8 we prove that the (Euclidean) conditioning $\kappa_{2}(V_{d})$ is bounded, and in Theorem 3.14 that the (Frobenius) conditioning $\kappa_{F}(V_{d})$ is linear. Both these results are provided with simple formulas that depend on $\rho$.
@@ 2 Ill-conditioning: the monomial basis
$\kappa_{2}\left(V_{d}\right)=\frac{\sigma_{1}\left(V_{d}\right)}{\sigma_{d}\left(V_{d}\right)}=\sqrt{\frac{\lambda_{1}\left(V_{d}^{\top}V_{d}\right)}{\lambda_{d}\left(V_{d}^{\top}V_{d}\right)}}$
of the Vandermonde matrix $V_{d}$ defined in (2) is exponential when the monomial basis $b_{j}(x):=x^{j-1}$ is chosen. As we shall see, this holds independently of the underlying collection of unisolvent segments. We build such a general result by first consdering the peculiar case of the class (C3), which yields very neat computations, and then show how our proof extends to other scenarios. To fix the ideas we start with a motivating example, whose features turn out to be quite general.
@@ 2.1 A motivating example
Consider $s_{i}=[0,i]$, for $i=1,\ldots,d$ and the monomial basis. The corresponding Vandermonde matrix is
@@
$\left[V_{d}^{\top}V_{d}\right]_{i,j}=\frac{1}{i\cdot j}\sum_{k=1}^{d}k^{i+j}=:\left[W_{d}\right]_{i,j}.$
@@
[…]
The motivating example is important for the following reason. Suppose that there exists $c>0$ such that $\sup_{i}\left(\beta_{i}-\alpha_{i}\right)\geq c$. When there exists $v>1$ independent of $d$ such that $\sup_{i,j}\max\{|\beta_{i}|,|\alpha_{j}|\}\geq v$, by direct inspection as above we have necessarily exponential growth for the maximal eigenvalue of $W_{d}=V_{d}^{\top}V_{d}$, with its minimal eigenvalue that cannot grow exponentially with respect to $d$. Hence, the latter gives a clear evidence of an exponential conditioning of $W_{d}$ and a fortiori of $V_{d}$ as a function of the size $d$, whenever $c>0$, $v>1$ are independent of the polynomial degree $d$. In what follows, we shall thus make a similar regularity assumption on $c$.
Now, observe that dividing by the size $d$ of the longest support $s_{d}$ one translates the problem in a fixed reference interval. By applying the change of variable $t=x/d$, one immediately sees that the $j$-th column of $V_{d}$ scales by the factor $1/d^{j}$ and, as a consequence, the $(i,j)$-th term of $W_{d}$ scales by $1/d^{i+j}$. This shows that the conditioning of $V_{d}$ depends on the length of the segments. In particular, even if it is easy to numerically check that the conditioning in this situation is still exponential, this example complicates the estimate on $\lambda_{1}$ given above and suggest that the reference interval I=[1,1]I=[-1,1] is somewhat a limiting case of the problem. This will guide our proof strategy.
@@ 2.1.1 A formula for the determinant for class (C3)
Before moving further, we extend [5, Eq. (24)], and obtain an easy formula for the determinant of $V_{d}$ for the class (C3) when the left endpoint is $\alpha_{i}=\alpha$. The special case $\alpha=0$ clearly applies to the motivating example.
@@
(3) $\det V_{d}=\frac{1}{d!}\left(\prod_{i=1}^{d}\left(\beta_{i}-\alpha\right)\right)\left(\prod_{1\leq j<k\leq d}\left(\beta_{k}-\beta_{j}\right)\right).$
Proof 2.2.
@@
The matrix appearing at the right hand side of the above equation is the Vandermonde matrix associated with the nodes $\{\beta_{1},\ldots,\beta_{d}\}$, whose determinant is well-known to be $\prod_{1\leq j<k\leq d}\left(\beta_{k}-\beta_{j}\right)$. The claim immediately now follows from Binet’s Theorem.
The formula (3) assumes the very neat form $\det(V_{d})=\prod_{i=1}^{d-1}i!$ under the assumptions of Section 2.1. This is readily seen plugging $\alpha=0$ and $\beta_{i}=i$ in (3). Lemma 2.1 also gives an alternative, more direct proof of the unisolvence of the class (C3) proposed in [4].
@@ 2.1.2 An application to the Fekete problem
(4) $\left|\det\left(\left[\frac{1}{|s_{i}|}\int_{s_{i}}b_{j}(x)\mathrm{d}x\right]_{i,j}\right)\right|$
@@
where we also removed the absolute value of (4) due to the positivity result given in Lemma 2.1. Since $\frac{1}{d!}$ is a constant term, $\det\widetilde{V}_{d}$ is maximixed by nodes $\{\beta_{1},\ldots,\beta_{d}\}$ that maximize the rightmost term of the above equality. Such nodes are well-known to be the Fekete points, which is the claim. Notice that the case $\alpha=\beta_{1}$ also follows by an immediate continuity argument.
@@ 2.2 A conditioning result for the reference interval
To extend the computations of Section 2.1, we start from the situation where the reference interval is I=[1,1]I=[-1,1]. This framework offers a neat handling due to the well-known features of Chebyshev polynomials.
@@
Let $e_{i}$ be the $i$-the element of the canonical basis of $\mathbb{R}^{d}$. Then, lower bounding the $2$-norm of the vector $V_{d}e_{i}$ by its first component, one has
$\left\|V_{d}e_{i}\right\|_{2}=\left\|\left[\frac{\beta_{i}^{j}-\alpha_{i}^{j}}{j}\right]_{j=1}^{d}\right\|_{2}\geq\left|\beta_{i}-\alpha_{i}\right|=\beta_{i}-\alpha_{i},$
whence
$\left\|V_{d}\right\|_{2}=\max_{\|u\|_{2}=1}\left\|V_{d}u\right\|_{2}\geq\max_{1\leq i\leq d}\left\|V_{d}e_{i}\right\|_{2}\geq\max_{1\leq i\leq d}\beta_{i}-\alpha_{i}\geq c,$
for a $c$ independent of the polynomial degree (hence of the matrix size) $d$. Now, in order to estimate from below the conditioning of $V_{d}$ in the Euclidean norm, we look for a proper estimation from below of $\left\|V_{d}^{-1}\right\|_{2}$. In particular, observe that for any $y\in\mathbb{R}^{d}$ one has
(5) $\left\|V_{d}^{-1}\right\|_{2}=\max_{\begin{subarray}{c}z\in\mathbb{R}^{d}\\ z\neq 0\end{subarray}}\frac{\|V_{d}^{-1}z\|_{2}}{\|z\|_{2}}\geq\frac{\|y\|_{2}}{\left\|V_{d}y\right\|_{2}}.$
It follows that, if we are able to select an appropriate $y\in\mathbb{R}^{d}$, we obtain a satisfactory estimate for $\left\|V_{d}^{-1}\right\|_{2}$. To this end, consider the Chebyshev polynomial $p_{d-1}^{*}(x):=\sum_{j=1}^{d}a_{j}x^{j-1}$ of degree d1d-1. This is a monic polynomial, so $a_{d}=1$, and it is well-known that its norm on [1,1][-1,1] is $\|p_{d-1}^{*}\|_{\infty}=2^{1-d}$. We evaluate the inverse of $V_{d}$ on the transformed vector $b=V_{d}a$. By construction, we have
$b_{i}=\int_{\alpha_{i}}^{\beta_{i}}p_{d-1}^{*}(z)\mathrm{d}z,$
whence, by the Hölder inequality,
$|b_{i}|=\left|\int_{\alpha_{i}}^{\beta_{i}}p_{d-1}^{*}(z)\mathrm{d}z\right|\leq(\beta_{i}-\alpha_{i})\|p_{d-1}^{*}\|_{\infty}\leq(\beta_{i}-\alpha_{i})2^{1-d}.$
@@
$\|b\|_{2}\leq\max_{1\leq i\leq d}(\beta_{i}-\alpha_{i})\frac{\sqrt{d}}{2^{d-1}}\leq 2\frac{\sqrt{d}}{2^{d-1}}=\frac{\sqrt{d}}{2^{d-2}}.$
By (5) we immediately deduce that $\left\|V_{d}^{-1}\right\|_{2}\geq 1/\|b\|_{2}\geq 2^{d-2}/\sqrt{d}$. Gathering all the results, we have
@@
Since $c$ does not depend on the degree $d$, the claim is proved.

### v2

# On the conditioning of polynomial histopolation

@@ Abstract
Histopolation is the approximation procedure that associates a degree d1d-1 polynomial $p_{d-1}\in\mathscr{P}_{d-1}(I)$ with a locally integrable function $f$ imposing that the integral (or, equivalently, the average) of $p$ coincides with that of $f$ on a collection of $d$ distinct segments $s_{i}$. In this work we discuss unisolvence and conditioning of the associated matrices, in an asymptotic linear algebra perspective, i.e., when the matrix-size $d$ tends to infinity. While the unisolvence is a rather sparse topic, the conditioning in the unisolvent setting has a uniform behavior: as for the case of standard Vandermonde matrix-sequences with real nodes, the conditioning is inherently exponential as a function of $d$ when the monomial basis is chosen. In contrast, for an appropriate selection of supports, the Chebyshev polynomials of second kind exhibit a bounded conditioning. A linear behavior is also observed in the Frobenius norm.
@@
Polynomial histopolation, conditioning, Chebyshev polynomials
@@ 1 Introduction
Let $I\subset\mathbb{R}$ be a compact interval and let $f:I\to\mathbb{R}$ be a locally integrable function. Histopolation is the approximation procedure that associates with $f$ a degree d1d-1 polynomial $p_{d-1}\in\mathscr{P}_{d-1}(I)$ such that
$\int_{\alpha_{i}}^{\beta_{i}}f(x)\mathrm{d}x=\int_{\alpha_{i}}^{\beta_{i}}p(x)\mathrm{d}x,$ (1)
for an appropriate collection of supports $s_{i}:=[\alpha_{i},\beta_{i}]$, $i=1,\ldots,d$. In the literature [7, 24], an equivalent formulation of (1) containing normalization by the measure $|s_{i}|$ of $s_{i}$ is sometimes adopted; in what follows, we mainly stick to the non-normalized case. The main consequence of this choice is that the measures of the supports must be bounded away from zero; that is, we assume that there exists $\varepsilon>0$ independent of $d$ such that $|s_{i}|\geq\varepsilon$ for each $i=1,\ldots,d$. This is a rather natural assumption: even in the normalized case, it avoids the introduction of technical concepts such as regular sets [21]. However, we will also discuss situations where the normalization factor is present, with particular focus to limiting cases where segments collapse to points.
In contrast to the univariate nodal interpolation procedure, where any collection of $d$ distinct points is unisolvent for polynomial interpolation in $\mathscr{P}_{d-1}(I)$, such a neat recipe does not exist for histopolation, where the identification of specific classes is a hard task. It is indeed very easy to construct examples of $d$ distinct segments that are not unisolvent for histopolation in $\mathscr{P}_{d-1}(I)$.
Example 1.1.
Consider the family of segments [i,i][-i,i], for $i=1,\ldots,d$ with $d\geq 1$. Any non-vanishing linear combination of polynomials of odd degree less than or equal to d1d-1 satisfies
@@
This is only one of the many differences between nodal interpolation and histopolation. Such a dissimilarity is mostly motivated by the fact that any interval $s_{i}$ requires two parameters to be described (e.g., its endpoints $\alpha_{i}$ and $\beta_{i}$, or the midpoint $\tau_{i}$ and the length $|s_{i}|$), giving rise to a generally underdetermined problem. As a direct consequence of this, unisolvence results are confined to specific classes of segments. Many of them were identified in [6], and are:
@@
    Figure 1: The segments $s_{1}$, $s_{2}$ and $s_{3}$ form a chain of intervals. They are examples of non-equidistributed segments in the class (C1).
@@
    segments with uniform arc-length. For some $\tau_{i}\in(0,\pi)$ and $0<\rho<\pi$, one has $\alpha_{i}=\cos(\tau_{i}+\rho)$ and $\beta_{i}=\cos(\tau_{i}-\rho)$. If the nodes $\tau_{i}$ are the Chebyshev nodes and $\rho=\pi/(2d)$, we refer to the corresponding collection as that of Chebyshev segments;
@@
Of course, there are families of segments that belong to more than one class. An example of this are the so-called Chebyshev segments, that we will study in depth in Section 3. Indeed, for the particular choice of the arc-length parameter $\rho=\pi/(2d)$ and the equidistribution of the nodes in the semicircle, $\tau_{i}=\frac{(i-1/2)\pi}{d}$ for $i=1,\ldots,d$, they can be framed both in class (C1) and class (C2). By exploiting the normalized version of (1), segments solving the Fekete problem have subsequently been recognized as an additional scenario [7]. In Section 2.3, we exhibit a new class of segments which is unisolvent for polynomial histopolation. We will denote this class by (C4).
@@
$\left[V_{d}\right]_{i,j}:=\int_{s_{i}}b_{j}(x)\mathrm{d}x.$ (2)
To stress the connection of the results with the number $d$ of segments, we consistently index the Vandermonde matrix $V_{d}$, as in (2). This quantity is also the size of the matrix, and is upshifted by $1$ with respect to the relative polynomial degree.
Example 1.2.
The non-unisolvence of the set of segments introduced in the above example is also readily seen by looking at the Vandermonde matrix computed with respect to the monomial basis $b_{j}(x)=x^{j-1}$. For any even $j$, one has
@@
It follows that any even row of $V_{d}$ is equal to the zero vector. Since the invertibility of the Vandermonde matrix does not depend on the polynomial basis chosen, one deduces that such a set is not unisolvent for histopolation in $\mathscr{P}_{d-1}(I)$.
The problem of computing the features of a Vandermonde matrix, in particular concerning its conditioning and hence the identification of ill-conditioned bases, is a cornerstone of approximation theory [23]; of course, so is also the opposite problem. In the framework of nodal interpolation the problem was extensively studied, and several results concerning optimality [14, 15, 18], bad conditioning [3, 8], displacement structures [4], as well as their applications [19], have been obtained, endorsing the design of efficient algorithms. The majority of these results can be grouped into two clusters: results concerning the monomial basis, which in general offers unstable performances, and results concerning the Chebyshev (of first kind) basis, which can be used in combination with Chebyshev nodes to reverse such a situation.
Outline of the paper and main results. In contrast to nodal interpolation, very little is known about the conditioning of the Vandermonde matrix (2) associated with the histopolation problem, and about its asymptotics as a function of the polynomial degree $d$. The principal aim of this work is to fill this gap.
To do so, we dedicate Section 2 to ill-conditioning. We recognize that the (Euclidean) conditioning $\kappa_{2}$ of the matrix $V_{d}$ grows exponentially with $d$ when the monomial basis is chosen. This is proved first in the reference interval I=[1,1]I=[-1,1] and then extended to any collection of segments by Theorem 2.11. This feature recalls the nodal case, but is not a consequence of the mean value theorem, as one could expect. To broaden the range of applicability of the result, we also identify a new class of unisolvent segments, and use it to produce numerical evidences.
To restore hope in histopolation, in Section 3 we present results concerning well-conditioning. Indeed, we show that Chebyshev polynomials of the second kind, in combination with segments in the class (C2), give rise to a column-orthogonal matrix. The diagonal terms of $V_{d}^{\top}V_{d}$ are then explicitly determined in Proposition 3.1. As a consequence, in Theorem 3.13 we prove that the (Euclidean) conditioning $\kappa_{2}(V_{d})$ is bounded, and in Theorem 3.20 that the (Frobenius) conditioning $\kappa_{F}(V_{d})$ is linear. Both these results are provided with simple formulas that depend on $\rho$.
@@ 2 Ill-conditioning: the monomial basis
$\kappa_{2}\left(V_{d}\right)=\|V_{d}\|_{2}\|V_{d}^{-1}\|_{2}=\frac{\sigma_{1}\left(V_{d}\right)}{\sigma_{d}\left(V_{d}\right)}=\sqrt{\frac{\lambda_{1}\left(V_{d}^{\top}V_{d}\right)}{\lambda_{d}\left(V_{d}^{\top}V_{d}\right)}}$
of the Vandermonde matrix $V_{d}$ defined in (2) is exponential when the monomial basis $b_{j}(x):=x^{j-1}$ is chosen. In the above formula, as customary, $\sigma_{j}(V_{d})$ is the $j$-th singular value $\sigma_{j}^{2}(V_{d})=\lambda_{j}\left(V_{d}^{\top}V_{d}\right)$ of $V_{d}$, and the corresponding sequence
$\sigma_{1}(V_{d})\geq\sigma_{2}(V_{d})\geq\cdots\geq\sigma_{d}(V_{d})$
is non-increasingly ordered with $j=1,\ldots,d$. We will frequently set $W_{d}:=V_{d}^{\top}V_{d}$, and use both the definition and the characterization of $\kappa_{2}(V_{d})$ given above.
As we shall see, the exponential conditioning is intrinsic to the monomial basis, and holds independently of the underlying collection of unisolvent segments. We build such a general result by steps, confining first ourselves in a reference interval. To fix the ideas we start with a motivating example concerning the class (C3), which yields very neat computations, and whose features turn out to be quite general.
@@ 2.1 A motivating example
Consider $s_{i}=[0,i]$, for $i=1,\ldots,d$, and the monomial basis. The corresponding Vandermonde matrix is
@@
$\left[W_{d}\right]_{i,j}:=\left[V_{d}^{\top}V_{d}\right]_{i,j}=\frac{1}{i\cdot j}\sum_{k=1}^{d}k^{i+j}.$
@@
Remark 2.1.
Suppose that there exists $c>0$ such that $\sup_{i}\left|\beta_{i}-\alpha_{i}\right|\geq c$. When there exists $v>1$ independent of $d$ such that $\sup_{i,j}\max\{|\beta_{i}|,|\alpha_{j}|\}\geq v$, by direct inspection as above we have necessarily exponential growth for the maximal eigenvalue of $W_{d}=V_{d}^{\top}V_{d}$, with its minimal eigenvalue that cannot grow exponentially with respect to $d$. Hence, the latter gives a clear evidence of an exponential conditioning of $W_{d}$ and a fortiori of $V_{d}$ as a function of the size $d$, whenever $c>0$, $v>1$ are independent of the polynomial degree $d$. In what follows, we shall thus make a similar regularity assumption on $c$, and discuss when it can be weakened.
Observe that dividing by the size $|s_{d}|=d$ of the longest support $s_{d}$ one translates the problem into a fixed reference interval. Applying the change of variable $t=x/d$, one immediately sees that the $j$-th column of $V_{d}$ scales by the factor $1/d^{j}$ and, as a consequence, the $(i,j)$-th term of $W_{d}$ scales by $1/d^{i+j}$. This shows that the conditioning of $V_{d}$ depends on the length of the segments and, even if it is easy to numerically check that the conditioning in this situation is still exponential, this complicates the estimate on $\lambda_{1}$ given above, suggesting that the reference interval I=[1,1]I=[-1,1] is somewhat a limiting case of the problem. This motivates the particular emphasis we put on this situation.
@@ 2.1.1 A formula for the determinant for class (C3)
The collection of segments used in Section 2.1 is an instance of class (C3). We now extend [7, Eq. (24)] and obtain an easy formula for the determinant of $V_{d}$ for the class (C3) when the left endpoint is $\alpha_{i}=\alpha$. The special case $\alpha=0$ clearly applies to the motivating example.
@@
$\det V_{d}=\frac{1}{d!}\left(\prod_{i=1}^{d}\left(\beta_{i}-\alpha\right)\right)\left(\prod_{1\leq j<k\leq d}\left(\beta_{k}-\beta_{j}\right)\right).$ (3)
Proof.
@@
The matrix appearing at the right hand side of the above equation is the Vandermonde matrix associated with the nodes $\{\beta_{1},\ldots,\beta_{d}\}$, whose determinant is well-known to be $\prod_{1\leq j<k\leq d}\left(\beta_{k}-\beta_{j}\right)$. The claim now follows immediately from Binet’s Theorem. ∎
The formula (3) assumes the very neat form $\det(V_{d})=\prod_{i=1}^{d-1}i!$ under the assumptions of Section 2.1. This is readily seen plugging $\alpha=0$ and $\beta_{i}=i$ in (3). Lemma 2.2 also gives an alternative, more direct proof of the unisolvence of the class (C3) proposed in [6]. Obviously, one may retrace the same proof to obtain an analogous formula by fixing the right endpoint $\beta_{i}=\beta$.
Corollary 2.3.
Let $s_{i}=[\alpha_{i},\beta]$, with $\alpha_{i}<\alpha_{j}<\beta$ for $1\leq i<j\leq d$. The determinant of $V_{d}$ is positive, and
$\det V_{d}=\frac{1}{d!}\left(\prod_{i=1}^{d}\left(\beta-\alpha_{i}\right)\right)\left(\prod_{1\leq j<k\leq d}\left(\alpha_{k}-\alpha_{j}\right)\right).$
@@ 2.1.2 An application to the Fekete problem
$\left|\det\left(\left[\frac{1}{|s_{i}|}\int_{s_{i}}b_{j}(x)\mathrm{d}x\right]_{i,j}\right)\right|$ (4)
@@
where we also removed the absolute value of (4) due to the positivity result given in Lemma 2.2. Since $\frac{1}{d!}$ is a constant term, $\det\widetilde{V}_{d}$ is maximized by nodes $\{\beta_{1},\ldots,\beta_{d}\}$ that maximize the rightmost term of the above equality. Such nodes are well-known to be the Fekete points, which is the claim. Notice that the case $\alpha=\beta_{1}$, which represents the first segment shrinking to the point $\alpha$, also follows by an immediate continuity argument. ∎
@@ 2.2 A conditioning result for the reference interval
To extend the computations of Section 2.1, we start from the situation where the reference interval is I=[1,1]I=[-1,1]. This framework offers a neat handling due to the well-known features of Chebyshev polynomials [20].
@@
Since all such terms are positive, we can trivially lower bound $\|V_{d}\|_{2}$ with any of the terms of $V_{d}$. Hence
$\|V_{d}\|_{2}\geq\max_{1\leq i,j\leq d}[V_{d}]_{i,j}\geq\max_{1\leq i\leq d}[V_{d}]_{i,1}=\max_{1\leq i\leq d}(\beta_{i}-\alpha_{i})\geq c,$
for a $c$ independent of the polynomial degree (hence of the matrix size) $d$. Now, in order to estimate from below the conditioning of $V_{d}$ in the Euclidean norm, we look for a proper estimation from below of $\|V_{d}^{-1}\|_{2}$. In particular, observe that for any non-zero $y\in\mathbb{R}^{d}$ one has
$\left\|V_{d}^{-1}\right\|_{2}=\max_{\begin{subarray}{c}z\in\mathbb{R}^{d}\\ z\neq 0\end{subarray}}\frac{\|V_{d}^{-1}z\|_{2}}{\|z\|_{2}}\geq\frac{\|y\|_{2}}{\left\|V_{d}y\right\|_{2}}.$ (5)
It follows that, if we are able to select an appropriate $y\in\mathbb{R}^{d}$, we obtain a satisfactory estimate for $\|V_{d}^{-1}\|_{2}$. To this end, we take $y^{\top}=(a_{1},\ldots,a_{d})$ as the vector of coefficients of the normalized Chebyshev polynomial of the first kind and degree d1d-1,
$p_{d-1}^{*}(x):=\frac{T_{d-1}(x)}{2^{d-1}}.$
Due to the normalization, $p_{d-1}^{*}(x)=\sum_{j=1}^{d}a_{j}x^{j-1}$ is a monic polynomial, so $a_{d}=1$, and this implies that $\|y\|_{2}\geq|a_{d}|=1$. Further, by construction, one has
$V_{d}y]_{i}=\int_{\alpha_{i}}^{\beta_{i}}p^{*}_{d-1}(x)\mathrm{d}x.$
Since $\|p_{d-1}^{*}\|_{\infty}=2^{1-d}$, by the Hölder inequality one immediately gets
$\left|[V_{d}y]_{i}\right|=\left|\int_{\alpha_{i}}^{\beta_{i}}p_{d-1}^{*}(x)\mathrm{d}x\right|\leq(\beta_{i}-\alpha_{i})\|p_{d-1}^{*}\|_{\infty}\leq(\beta_{i}-\alpha_{i})2^{1-d}.$
@@
$\|V_{d}y\|_{2}\leq\max_{1\leq i\leq d}(\beta_{i}-\alpha_{i})\frac{\sqrt{d}}{2^{d-1}}\leq 2\frac{\sqrt{d}}{2^{d-1}}=\frac{\sqrt{d}}{2^{d-2}}.$
We may hence estimate (5) from below by $\|V_{d}^{-1}\|_{2}\geq 1/\|V_{d}y\|_{2}\geq 2^{d-2}/\sqrt{d}$. Gathering all the results, we have
@@
Since $c$ does not depend on the degree $d$, the claim is proved. ∎
@@
2.2.1 The normalized case
In the determination of Fekete segments, the presence of a normalization factor has proven to play an important role [7], giving a hybrid situation between the limit where all segments shrink to points and the technical request $\sup_{i}(\beta_{i}-\alpha_{i})\geq c>0$. From the perspective of the conditioning of the associated Vandermonde matrix
$\widetilde{V}_{d}]_{i,j}:=\frac{1}{\beta_{i}-\alpha_{i}}\int_{\alpha_{i}}^{\beta_{i}}x^{j-1}\mathrm{d}x$
this does not mark a relevant difference, as the next result shows.
Corollary 2.6.
Let $1\leq\alpha_{i}<\beta_{i}\leq 1$ for $i=1,\ldots,d$, and assume that the corresponding collection of segments is unisolvent for histopolation in $\mathscr{P}_{d-1}(I)$. Then
$\kappa_{2}\left(\widetilde{V}_{d}\right)\geq\frac{2^{d-1}}{\sqrt{d}}.$
Proof.
Following exactly the same lines of the proof of Theorem 2.5, we have
$\|\widetilde{V}_{d}\|_{2}\geq\max_{1\leq i,j\leq d}[\widetilde{V}_{d}]_{i,j}\geq\max_{1\leq i\leq d}[\widetilde{V}_{d}]_{i,1}=\max_{1\leq i\leq d}\frac{\beta_{i}-\alpha_{i}}{\beta_{i}-\alpha_{i}}=1.$
Likewise, the estimate using normalized Chebyshev polynomials gives
$\left|[\widetilde{V}_{d}y]_{i}\right|=\left|\frac{1}{\beta_{i}-\alpha_{i}}\int_{\alpha_{i}}^{\beta_{i}}p_{d-1}^{*}(x)\mathrm{d}x\right|\leq\frac{\beta_{i}-\alpha_{i}}{\beta_{i}-\alpha_{i}}\|p_{d-1}^{*}\|_{\infty}\leq 2^{1-d}.$
Hence, we conclude that $\kappa_{2}\left(\widetilde{V}_{d}\right)\geq\frac{2^{d-1}}{\sqrt{d}}$, improving the above result by the factor $c/2$. ∎
The subsequent extension of Theorem 2.5 to generic intervals will immediately carry over to the normalized case. Note that Corollary 2.6 does not require anymore the assumption $\sup_{i}(\beta_{i}-\alpha_{i})\geq c>0$, and is in fact blind to $c$. In this case, the histopolation problem can be properly defined on regular sets [21], and involves both averages and point evaluations.